OpenStudy (anonymous):

i need help... give an equation of a parabola that is wider than f(x) and has a vertex that is lower than the vertex of f(x)

OpenStudy (anonymous):

hello this is a basic graph, and i recommend memorizing them :) and then you can use your transformations from the basic functions

OpenStudy (anonymous):

do you know what the parabola would look like f(x)=x^2

OpenStudy (anonymous):

vertex would be at (0,0) correct?

OpenStudy (anonymous):

no . . . ?

OpenStudy (anonymous):

ok ill show you

OpenStudy (anonymous):

did they give any pictures to the problem?

OpenStudy (anonymous):

no they didnt :( it just says to give an equation

OpenStudy (anonymous):

ok thats fine, the function f(x)= x^2 will look like this

OpenStudy (anonymous):

do you know your transformations? how to move this function around and transform it

OpenStudy (anonymous):

you can make (x^2) wider by multiplying by a fraction 1/2(x^2)

OpenStudy (anonymous):

isnt it like 1,3,5 ??? right ??? i think my teacher show me once before but im not sure

OpenStudy (anonymous):

and the vertex is at (0,0) so to move it down we move down the y axis (0,-1) for example

OpenStudy (anonymous):

if you multiply the (x^2) by a whole number the parabola will get more narrow

OpenStudy (anonymous):

okay im kinda understanding now

OpenStudy (anonymous):

so if you have f(x)=x^2 and you want to make it lower and wider you will have f(x)=1/2(x^2)-1

OpenStudy (anonymous):

you can exchange the fraction and the -1 with any other fraction or negative y value

OpenStudy (anonymous):

if you want to move the (x^2) along the x axis it would look like ((x^2)-1) would move the vertex to the right one spot and ((x^2)+1) would move the vertex left one spot

OpenStudy (anonymous):

but they aren't asking to move it left or right :0

OpenStudy (anonymous):

my recommendation is to look at your basic functions, and memorize them!

OpenStudy (anonymous):

so my answer would be f(x)=1/2(x^2)-1 . . . . ????

OpenStudy (anonymous):

yes'm or any variation of that

OpenStudy (anonymous):

it could be \[\frac{ 3 }{ 8 }(x^{2})-8\]

OpenStudy (anonymous):

the fraction makes it wider, and the -8 moves it down the y axis :)

OpenStudy (anonymous):

i do really need to look over my basic functions and thanks for the help! :)

OpenStudy (anonymous):

no prob have a good one hit me up if u have any other q's peace